OpenStudy (anonymous):

Part II: Create Your Own Conversion Problem Now it’s time for you to create your own problem that involves converting three different equations of lines to different forms. First, write three (3) equations of lines in the following forms: a) point-slope form, b) slope-intercept form, and c) standard form. Next, choose another form to write each equation in, and then show all the steps required to convert each equation to the other form.

OpenStudy (tacotime):

Do you know what point slope, slope intercept, and standard form are?

OpenStudy (anonymous):

yes i can put the ones i whant to convert in here if you want

OpenStudy (tacotime):

I guess. I don't understand what you need help on. Write all of your equations then convert them to some other form.

OpenStudy (anonymous):

yea i need help with converting them and showing all the steps.

OpenStudy (anonymous):

y=-1/2x + 4 to point slope form 8x + 2y =6 to slope intercept form y - 2 = 4 (x + 1) to standard form

OpenStudy (tacotime):

ok slope intercept is y=mx+b point slope is y-y1=m(x-x1) standard is mx+y=b

OpenStudy (anonymous):

ok but i have to convert the numbers over to

OpenStudy (tacotime):

To convert y = 1/2x+4 to point slope form first pick a random number for x. I'll chose 0 to make things easy. Plug it into the formula to get y = 1/2 * 0 + 4 thus y = 4 now plug y=4 into the original formula 4 = 1/2 x +4 x = 2 Now use these in the point slope form along with m = -1/2 y - 4 = -1/2 (x - 2)

OpenStudy (anonymous):

so m =-1/2 y - 4 = -1/2 (x -2) is the answer for y=-1/2x + 4

OpenStudy (tacotime):

Yeah its just y - 4 = -1/2 (x -2) because that's point slope form

OpenStudy (anonymous):

ok thanks can you help me with the others

OpenStudy (tacotime):

8x + 2y =6 to slope intercept form For this one you want to isolate y 2y = 6 - 8x y = 3 - 4x y = -4x + 3 that's it

OpenStudy (anonymous):

2y = 6 - 8x y = 3 - 4x y = -4x + 3 is the answer and the work

OpenStudy (anonymous):

ok thanks, so much this is helping me alot

OpenStudy (anonymous):

y - 2 = 4 (x + 1) to standard form?

OpenStudy (tacotime):

y - 2 = 4 (x + 1) to standard form First distribute 4 y - 2 = 4x + 1 then isolate y y = 4x + 2 y-4x = 2 -4x + y = 2